2 The coffee industry claims that the average U.S. adult drinks 1.7 cups of coffee per day. a) State the Hypothesis to show the average cups of coffee per day is different than 1.7. b) Choose a level of a. Use a = 0.01 for this problem. c) To test the hypothesis, a random sample of 34 adults was selected and their average coffee consumption was recorded. The data appear in the Coffee worksheet of the HW1 data workbook on Moodle. Collect data and calculate necessary statistics to test the hypothesis. d) Sketch the sampling distribution. Include the critical value and test statistic e) Draw a conclusion and report that in the problem context.

0
5
1
0
4
5
6
4
2
4
5
4
2
1
3
2
7
1
3
0
3
1
4
0
7
2
5
5
4
3
1
1
0
0
Answer #1

Ans:

a)

Ho : μ = 1.7

Ha : μ 1.7

b)

df=34-1=33

alpha=0.01

critical t value=tinv(0.01,33)=+/2.733

c)

Sample mean=2.794

sample standard deviation=2.115

n=34

d)

Test statistic:

t=(2.794-1.7)/(2.115/sqrt(34))

t=3.016

As, test statistic falls in rejection region,we reject the null hypothesis.

There is sufficient evidence to conclude that the average cups of coffee per day is different than 1.7.

0 几 3 43210